The perimeter of a rectangle is 104 inches. The length exceeds the width by 14 inches.

Answer by tm.hodgson(1)

P=2l + 2w
104=2(w+14)+2w
104=2w + 28 + 2w
104-28=4w
76=4w
19=w
104=2l +2(19)
104=2l +38
104-38=2l
66=2l
33=l